OpenStudy (anonymous):

what is the reference angle of 7pi/4?

OpenStudy (anonymous):

well\[0\le \frac{7\pi}{4} \le 2\pi\]so the reference angle must be itself if i'm not wrong

OpenStudy (anonymous):

what do u think?

OpenStudy (anonymous):

wait...can you explain it a different way? I'm having a hard time understanding.

OpenStudy (anonymous):